Документ взят из кэша поисковой машины. Адрес оригинального документа : http://www.stsci.edu/smov/notes/0218smov.html
Дата изменения: Tue Mar 11 23:58:42 1997
Дата индексирования: Thu Feb 27 23:53:31 2014
Кодировка:

Поисковые слова: http www.badastronomy.com phpbb index.php
SMOV Status & Planning Meeting Minutes 02/18/97

SMOV Status & Planning Meeting Minutes

Meeting Date: 02/18/97


STATUS REPORTS

1.  SM & Observatory Status

        - EVA 5 completed successfully.  
        Concern over slow RWA2 speeds was alleviated 
        with a test that showed proper operation.

        - Final reboost results in a 335 x 320 n.m. orbit.
        As a result a STIS basefile parameter for orbital 
        period will need to be adjusted.

        - SSR indicates some SEUs in one internal control bus
        during SAA passages that, if occurring during a data 
        transer to/from memory, could result in loss of 1 block
        of data.  (~250kb @ 1Mbs data rate).  Since not normally
        recording/dumping during SAA, it is not considered
        a serious problem.

        - Other spacecraft status is nominal.
 
2.  WFPC2 Status - nominal

3.  NICMOS Status 
        - eng. status is nominal
        - science status reports an unexpectedly large shift
          in the position of the coronagraphic hole in Camera 2
          and some vignetting in camera 3.  Believed to be due
          to dewar/optical_bench misalignment possibly caused
          by gravity release and heat soaking while in the SIPE.
          Not believed to affect our ability to focus
          during SMOV.  May improve with temperature change of
          dewar outer shell.  Proposal to re-run a portion of the
          NICMOS functional test is being considered.

4. STIS Status

        - nominal
        - CCD TEC at -66 Deg. C

5.  FGS status

        - nominal

6.  FOC/COSTAR status

        - nominal

7.  PLANNING REPORTS

        1.  Status of Upcoming SMSs

                - HS051 (H&S SMS) - preliminary sms is complete.  May get
                        another delivery from GSFC after Release,
                        but not a problem because 11-day ehpem not 
                        due until 6 am. Wed. 2/19.

                The goal is to run all of the following SMSs by Sunday for input
                        to commanding review in order to get a jump on TDRS
                        requests in forecast period.

                - SMS053 (1st SMOV Science SMS) - scheduled to be built with
                        11-day ephem and may be rebuilt with 10-week
                        ephem due at 1pm Wed., 2/19.

                - SMS055 - replanned test calendar complete
                         - TDRSS contacts being negotiated today.
 
                - SMS058 - replanned test calendar complete     
                         - TDRSS contacts being negotiated today.
                         - need contingency SMS058 in case WFPC2 contamination
                           is seen in SMS055.

                - SMS062 - replanned test calendar complete 

                - SMS065- replanned test calendar complete

                - SMS069- replanned test calendar complete

                - SMS072- replanned test calendar complete

                Merle has provided an SMS/PASS processing schedule for
deliveries
                though early March.
 

       2.  Replans - none

8.  OPERATIONAL ISSUES

        None



biagetti@stsci.edu

Go Back Go back to the Meeting Minutes Page